23 six less 5 times a number is subtracted from 3 times the number. If the result is 12, find the number.
Answer by ptaylor(1894)

SEE PROBLEM 69263

Let x= the number
Then five times the number is 5x
Six less than 5 times the number is 5x-6
Three times the number is 3x
6 less than 5 times a number, subtracted from 3 times the number is 3x-(5x-6) and this equals 12, we are told. So, our equation to solve is:

3x-(5x-6)=12 get rid of parens (to do that we have to multiply (5x-6) by -1)
3x-5x+6=12 subtract 6 from both sides
3x-5x+6-6=12-6 collect like terms
-2x=6 divide both sides by -2
x=-3
CK
3x-(5x-6)=12
3(-3)-(5(-3)-6)=12
-9-(-15-6)=12
-9-(-21)=12
-9+21=12
12=12
Hope this helps-----ptaylor